# Multiclass Local Calibration With the Jensen-Shannon Distance

### v1

@@ 5 Evaluating Multiclass Calibration
Let $f$ be a probabilistic classifier, and let $D$ denotes an evaluation dataset. Let $\beta:\Delta^{C}\rightarrow\{1,\ldots,m_{B}\}$ be a deterministic binning function that partitions the probability simplex $\Delta^{C}$ into $m_{B}$ disjoint bins $\{B_{b}\}_{b=1}^{m_{B}}$ and let $\varphi:[0,1]\times[0,1]\rightarrow\mathbb{R}_{\geq 0}$ be a scalar comparator that measures discrepancy between an empirical frequency $\text{freq}_{b,c}$ and a predicted confidence $\text{conf}_{b,c}$ that is Lipschitz in both arguments with constant $L_{\varphi}$. The general multiclass bin-based calibration error metric is obtained as:
$\mathcal{E}(D;\varphi;\beta)=\sum_{b=1}^{m_{B}}w_{b}\sum_{c=1}^{C}\pi_{c}\cdot\varphi(\text{freq}_{b,c},\text{conf}_{b,c}).$ (4)
where $w_{b}$ are deterministic bin weights and $\pi_{c}$ are deterministic class weights.
This general definition of calibration metrics provides a flexible framework that subsumes all calibration measures that leverage binning, including $ECE$ and $MECE$. Interestingly, we can define a probabilistic bound for the value of any calibration metric satisfying definition 2 under local calibration.
@@
Let $\mathcal{E}(D;\varphi;\beta)$ denote a calibration metric according to definition 2. If a classifier $f$ satisfies local calibration for error $\varepsilon$, then, for any $\delta\in[0,1]$, with at least probability $1-\delta$ it holds:
$\mathcal{E}(D;\varphi;\beta)\leq L_{\varphi}\cdot\Bigg[\varepsilon+\sum_{b=1}^{m_{B}}w_{b}\sqrt{\frac{\log(2Cm_{B}/\delta)}{2|\Psi(b,\mathcal{Y})|}}\Bigg],$ (5)
where $\Psi(\cdot;\cdot)$ selects a bin based on the index $b$ and the set of labels $\mathcal{Y}$.
We provide the proof in the Appendix B. Notice that Theorem 2 establishes an upper bound on $\mathcal{E}(D;\varphi;\beta)$ in terms of $\varepsilon$, implying that a high value of the metric necessarily indicates poor local calibration. However, the converse does not hold in general, as binning-based metrics are known to suffer from cancellation effects, whereby opposing errors may offset each other and obscure miscalibration.
@@ 6 Evaluating Calibration in Neural Networks
In the following section, we focus on Neural Network classifiers. Exploiting their structure and the $L$-Lipschitz continuity property with respect to the norm $\|\cdot\|_{1}$ [4], we refine the general results presented in the previous sections and investigate further the properties of Local Calibration Error. Practically speaking, achieving perfect local calibration is unfeasible. In the ideal case where $\varepsilon=0$, the distribution of points would be both perfectly centered around $\mathbf{x}_{i}$ and fully representative of the local data distribution. However, such conditions are never met in finite samples. Additionally, definition 1 does not take into account potential disparities in the miscalibration error of classes but only bounds the total sum of errors. For these two reasons, we relax definition 1 by bounding the admissible error in the predicted probability for class $c$ by the maximum change in the output probabilities that can occur when moving within a ball of radius $\rho$ in the feature space. Thus, decreasing $\rho$ bounds the class-wise deviation a model can tolerate while approaching perfect local calibration. Note that our definition operates under the assumption that the kernel estimator refers to the Neural Network learned feature-representations learned as in [15].
@@
Let $f$ be a probabilistic classifier composed of a feature extractor $\phi:\mathcal{X}\to\mathcal{F}$ and a final classification layer $g:\mathcal{F}\to\Delta^{C}$. Let $D$ be an evaluation set. In addition, $\phi(\mathbf{x}_{i})$ is assumed Lipschitz-continuous with respect to the softmax with constant $L>0$. We say that the model $f$ is $\rho$-perfectly locally calibrated if for every instance $\mathbf{x}_{i}\in D$ and for every class $c\in\{1,\ldots,C\}$, the absolute calibration error is bounded as follows:
@@
$\limsup_{\rho\to 0}\mathcal{E}(D;\varphi;\beta)\leq\sum_{b=1}^{m_{B}}w_{b}\sqrt{\frac{\log(2Cm_{B}/\delta)}{2|\Psi(b,\mathcal{Y})|}}$ (7)
Proof can be found in Appendix. This result underscores that the smaller the value of $\rho$ for which a model exhibits perfect local calibration, the more the overall error is dominated by the stochastic component.
Our theoretical analysis on the role of local calibration in shaping the behavior of calibration metrics will next address the multiclass version of $LCE$. Its inclusion is essential as it is the only metric that directly quantifies the degree of local calibration of a probabilistic classifier. More precisely, we show that, unlike global calibration measures, LCE relies on kernel regression estimates, which introduces a bias–variance trade-off on top of the inherent calibration error. Roughly speaking, the bound we derive decomposes into three interpretable components: (i) a calibration term $\varepsilon$ controlled by the local calibration property; (ii) a variance term growing as the kernel weights become concentrated on fewer neighbors; and (iii) a bias term which penalizes assigning large weights to distant samples.
@@
Let $f$ be a probabilistic classifier, where $f$ is composed of a feature extractor $\phi$ and a final classification layer $g$. Assume $f$ to be locally calibrated up to error $\varepsilon$ and let the LCE be computed on $\phi(\mathbf{x}_{i})$. Then, there exists $k\in[1/C,1]$ such that, for any $\delta\in[0,1]$, with at least probability $1-\delta$ the following holds:
$\mathrm{LCE}\leq+k\Bigg[\varepsilon+\overbrace{\frac{1}{n}\sum_{b=1}^{m_{B}}\sum_{i\in I_{b}}\sqrt{\frac{2\log\!\bigl(\tfrac{n}{\delta}\bigr)}{n_{i}^{\mathrm{eff}}}}}^{\text{variance term}}+\underbrace{\frac{L}{n}\sum_{b=1}^{m_{B}}\mathbb{E}\!\Big[\sum_{i\in I_{b}}\sum_{j\in I_{b}}w_{i,j}\|\phi(\mathbf{x}_{i})-\phi(\mathbf{x}_{j})\|_{1}\Big]\Bigg]}_{\text{bias term}}$
@@
We provide the proof in Appendix. This decomposition highlights the fundamental bias–variance trade-off induced by kernel smoothing: tighter kernels reduce bias at the expense of increased variance, while broader kernels reduce variance but incur larger bias. Therefore, estimating the order of magnitude of these terms clarifies the state of local calibration for a model.
@@ 7 Improving Local Calibration in Practice
We now focus on improving the local calibration of Neural Networks in practice. We first introduce Local Calibration Networks (LCN), i.e., a two-component neural network architecture, designed to produce representations that exhibit improved local calibration.
Figure 1(a) shows LCNs structure. While standard post hoc calibrators consider the features representation as fixed, one component of LCNs produces a new reduced-dimensionality feature representation $\phi^{\prime}(\mathbf{x}_{i})$ and the other component parametrizes new logits $\mathbf{l}^{\prime}$. These two components can be trained to minimize the misalignment between the probabilistic outputs and the local estimates of the class distribution computed from $\phi^{\prime}(\mathbf{x}_{i})$. To obtain these estimates, one can restort to kernel-based methods, e.g., Nadaraya-Watson estimators [17, 27], which require specifying a bandwidth hyper-parameter $\gamma$. More precisely, LCNs minimize the following loss:
$\mathcal{L}_{\text{lcl}}(\mathbf{x}_{i},\mathbf{y}_{i})=\frac{1}{n}\sum_{i=1}^{n}\overbrace{\text{d}_{\mathrm{JSD}}\left(\mathbf{\hat{p}}_{i},\hat{\theta}(\mathbf{y}_{i}\mid\phi^{\prime}(\mathbf{x}_{i})\right)}^{\text{Alignment term}}+\lambda\cdot\underbrace{\mathcal{L}_{\text{ce}}\left(\mathbf{y}_{i},\hat{\theta}(\mathbf{y}_{i}\mid\phi^{\prime}(\mathbf{x}_{i})\right)}_{\text{Similarity term}},$ (8)
@@
On the one hand, the alignment term leverages the Jensen-Shannon distance between the model’s predicted probability vector and the kernel estimates for each instance in the training batch. We prove that, for a consistent estimator, the alignment term asymptotically converges to the divergence between the model prediction distribution and the true label distribution.
@@
For a probabilistic classifier and a consistent (in the MSE sense) estimator of point-wise conditional probabilities, the average Jensen-Shannon distance between the model confidences and the kernel estimates converges to the one computed using the true distribution $\mathbf{p}$, i.e.,:
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}\text{d}_{\mathrm{JSD}}\left(\mathbf{\hat{p}}_{i},\hat{\theta}(\mathbf{y}\mid\mathbf{x}_{i}\right)=\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}\text{d}_{\mathrm{JSD}}\left(\mathbf{\hat{p}}_{i},\mathbf{p}_{i}\right).$ (9)
We provide the proof in the Appendix. Notably, the alignment term allows to learn probabilistic outputs that match the empirical neighbourhood frequencies.
On the other hand, the similarity term leverages the categorical cross entropy between the ground truth and the kernel estimates. Intuitively, this term encourages points with the same label to be attracted to nearby neighbours, but because the kernel is local, distant points of the same class exert little influence on each other. As a result, points that share fine-grained similarities are placed closer together, while more distinct variants remain further apart, but within the same class cluster. We provide a visual intuition of this behavior in Figure 1(b). However, the similarity term could still collapse class representations if not regularized or if the kernel bandwidth is too wide. Hence, we introduce another regularization hyper-parameter $\lambda$ (to be fine-tuned) to prevent this behaviour.
@@ 8 Experimental Evaluation
  • Q1: Does our method match the performance of baselines in the global calibration metric?
  • Q2: Does our proposal outperform baselines on local calibration metrics?
  • Q3: Does our approach affect predictions?
Datasets and Methods.
We evaluate our research questions over three multiclass datasets, i.e., cifar10, cifar100 [11], and tissuemnist from the MedMNIST collection [29]. We evaluate our approach, LCN, against publicly available calibration baselines, including Temperature Scaling (TS[5], Isotonic Regression (IR[31], and Platt Scaling (PS[20]. In addition, we compare with Dirichlet Calibration (DC[13], the current state-of-the-art for multiclass calibration. We train a separate model for each dataset: a ResNet-50 backbone for CIFAR-10 and TissueMNIST, and a deeper ResNet-152 for CIFAR-100. Hyperparameter settings are provided in the Appendix D. The code to reproduce our results can be found at https://github.com/Cesbar99/local-calibration/.
Metrics.
For Q1, we consider two global calibration metrics, i.e., $ECE$ and $ECCE$. For Q2, we compute two local calibration metrics, i.e., $LCE$ and $MLCE$. For Q3, we report two performance measures, i.e., accuracy ($Acc$) and Negative Log-Likelihood ($NLL$).
Experimental Setup.
To evaluate calibration for all methods, we separate the data into three disjoint parts: $(i)$ a training set (further split into training and validation); $(ii)$ a calibration set (with an internal calibration/validation split); and $(iii)$ a held-out test set used exclusively for evaluation. For CIFAR-10 and CIFAR-100, we use 45% of the original training data for training, 10% for validation, and 45% as test data, with the original test split serving as calibration data. For TissueMNIST, we use the pre-computed splits, but since calibration metrics computation requires larger sample sizes to be meaningful, we split the training data in half, yielding balanced training and test splits ($\approx 82.5k$ each). The remaining pre-computed test set ($\approx 40k$ instances) is used for calibration. For all the datasets, we split the calibration data into two sets, one to learn the calibration technique ($90\%$ of the calibration data) and one for validation (the remaining $10\%$ of the calibration data). Then, we $(i)$ train classifiers on the training set; $(ii)$ calibrate using the different methods on the calibration set; $(iii)$ compute the six metrics on the test set. We repeat this procedure using 5 different seeds and average results. Implementation details are provided in the Appendix D.
(a) Cifar10
(b) Cifar100
(c) TissueMNIST
@@
(a) Cifar10
(b) Cifar100
(c) TissueMNIST

### v2

@@ 5 EVALUATING MULTICLASS CALIBRATION
Let $\beta:\Delta^{C}\rightarrow\{1,\ldots,m_{B}\}$ be a deterministic binning function that partitions the probability simplex $\Delta^{C}$ into $m_{B}$ disjoint bins $\{B_{b}\}_{b=1}^{m_{B}}$ and let $\varphi:[0,1]\times[0,1]\rightarrow\mathbb{R}_{\geq 0}$ be a scalar comparator that measures discrepancy between an empirical frequency $\text{freq}_{b,c}$ and a predicted confidence $\text{conf}_{b,c}$ that is Lipschitz in both arguments with constant $L_{\varphi}$. The general multiclass bin-based calibration error metric is obtained as:
$\mathcal{E}(D;\varphi;\beta)=\sum_{b=1}^{m_{B}}\sum_{c=1}^{C}w_{b,c}\cdot\pi_{c}\cdot\varphi(\text{freq}_{b,c},\text{conf}_{b,c})$ (5)
where $w_{b,c}$ are possibly class dependent deterministic bin weights and $\pi_{c}$ are deterministic class weights.
This general definition of calibration metrics provides a flexible framework that subsumes all calibration measures that leverage binning, including $ECE$ and $MECE$. Interestingly, we can define a probabilistic bound for the value of any calibration metric satisfying Definition 2 under local calibration. See Appendix A for a detailed description of the assumptions.
@@
Let $\mathcal{E}(D;\varphi;\beta)$ be a calibration metric satisfying Definition 2, evaluated on a sample $\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{n}$. Let $f$ be locally calibrated with error $\varepsilon$ relative to kernel estimates $\hat{\theta}(\mathbf{y}_{i}\mid\mathbf{x}_{i})$ computed on an independent arbitrarily large auxiliary dataset. Conditioning on the evaluation inputs $\{\mathbf{x}_{i}\}_{i=1}^{n}$ and the corresponding consistent kernel estimates, for any $\delta\in(0,1)$, with probability at least $1-\delta$:
$\mathcal{E}(D;\varphi;\beta)\leq L_{\varphi}\sum_{b=1}^{m_{B}}\sum_{c=1}^{C}w_{b,c}\pi_{c}\Bigg(\sqrt{\frac{\log(\frac{2Cm_{B}}{\delta})}{2|\Psi(b,c)|}}\;+\;\varepsilon\Bigg)$ (6)
where $\Psi(\cdot,\cdot)$ selects a bin based on the index $b$ and possibly the label $c$.
We provide the proof in the Section B.2 together with a discussion of the result under limited cardinality of the auxiliary dataset leveraged for kernel estimates. We also show the result extends to cumulative metrics like $ECCE$ in Appendix C.1. The bound reveals a natural decomposition of calibration error into two components: a deterministic term proportional to the local calibration error $\varepsilon$, and a stochastic term that arises inevitably from finite-sample estimation of empirical frequencies within bins. Even when a model is perfectly locally calibrated ($\varepsilon=0$), the calibration metric remains subject to sampling fluctuations that vanish only as bin cardinalities grow. Notice that Theorem 2 establishes an upper bound on $\mathcal{E}(D;\varphi;\beta)$ in terms of $\varepsilon$, implying that a high value of the metric necessarily indicates poor local calibration. However, the converse does not hold in general, as binning-based metrics are known to suffer from cancellation effects, whereby opposing errors may offset each other and obscure miscalibration.
@@ 6 EVALUATING CALIBRATION IN NEURAL NETWORKS
In the following section, we focus on Neural Network classifiers. Exploiting their structure and the $L$-Lipschitz continuity property with respect to the norm $\|\cdot\|_{1}$ (Gao and Pavel, 2017), we refine the general results presented in the previous sections and investigate further the properties of Local Calibration Error. Practically speaking, achieving perfect local calibration is unfeasible. In the ideal case where $\varepsilon=0$, the distribution of points would be both perfectly centered around $\mathbf{x}_{i}$ and fully representative of the local data distribution. However, such conditions are never met in finite samples. Additionally, Definition 1 does not take into account potential disparities in the miscalibration error of classes but only bounds the total sum of errors. For these two reasons, we relax Definition 1 by bounding the admissible error in the predicted probability for class $c$ by the maximum change in the output probabilities that can occur when moving within a ball of radius $\rho$ in the feature space. Thus, decreasing $\rho$ bounds the class-wise deviation a model can tolerate while approaching perfect local calibration. Note that our definition operates under the assumption that the kernel estimator refers to the Neural Network learned feature-representations as in Luo et al. (2022).
@@
Let $f$ be composed of a feature extractor $\phi:\mathcal{X}\to\mathcal{F}$ and a classification layer $g:\mathcal{F}\to\Delta^{C}$. In addition, $\phi(\mathbf{x}_{i})$ is assumed Lipschitz-continuous with respect to the softmax with constant $L>0$. We say that $f$ is $\rho$-perfectly locally calibrated if for every instance $\mathbf{x}_{i}\in D$ and for every class $c\in\{1,\ldots,C\}$, the absolute calibration error is bounded as follows:
@@
$\limsup_{\rho\to 0}\mathcal{E}(D;\varphi;\beta)\leq L_{\varphi}\sum_{b=1}^{m_{B}}\sum_{c=1}^{C}w_{b,c}\pi_{c}\sqrt{\frac{\log(\frac{2Cm_{B}}{\delta})}{2|\Psi(b,c)|}}$ (8)
Proof is in Section B.3. This result underscores that the smaller the value of $\rho$ for which a model exhibits perfect local calibration, the more the overall error is dominated by the stochastic component.
Our theoretical analysis on the role of local calibration in shaping the behavior of calibration metrics will next address the multiclass version of $LCE$. Its inclusion is essential as it is the only metric that directly quantifies the degree of local calibration of a probabilistic classifier. More precisely, we show that, unlike global calibration measures, LCE relies on kernel regression estimates, which introduces a bias–variance trade-off on top of the inherent calibration error. Roughly speaking, the bound we derive decomposes into three interpretable components: (i) a calibration term $\varepsilon$ controlled by the local calibration property; (ii) a variance term growing as the kernel weights become concentrated on fewer neighbors; and (iii) a bias term which penalizes assigning large weights to distant samples. We discuss in detail assumptions in Appendix A.
@@
Let $f$ be composed of a feature extractor $\phi$ and classification layer $g$. Let $f$ be locally calibrated with error $\varepsilon$ relative to kernel estimates $\hat{\theta}(\mathbf{y}_{i}\mid\mathbf{x}_{i})$ computed on an independent arbitrarily large auxiliary dataset. Let the LCE be computed on $\phi(\mathbf{x}_{i})$. Then, conditioning on the evaluation inputs $\{\mathbf{x}_{i}\}_{i=1}^{n}$ and on the corresponding consistent kernel estimates, for any $\delta\in[0,1]$, with at least probability $1-\delta$:
$\mathrm{LCE}\leq\;\varepsilon\;+\;\overbrace{\frac{2}{n}\sum_{b=1}^{m_{B}}\sum_{i\in I_{b}}\sqrt{\frac{\log\!\bigl(\frac{2nC}{\delta}\bigr)}{2n_{i}^{\mathrm{eff}}}}}^{\text{variance term}}$
$\>+\>$ $\underbrace{\frac{L}{n}\sum_{b=1}^{m_{B}}\mathbb{E}\!\Big[\sum_{i\in I_{b}}\sum_{j\in I_{b}}w_{i,j}\|\phi(\mathbf{x}_{j})-\phi(\mathbf{x}_{i})\|_{1}\Big]}_{\text{bias term}}$
@@
We provide the proof in Section B.4. This decomposition highlights the fundamental bias–variance trade-off induced by kernel smoothing: tighter kernels reduce bias at the expense of increased variance, while broader kernels reduce variance but incur larger bias. Therefore, estimating the order of magnitude of these terms clarifies the state of local calibration for a model.
@@ 7 IMPROVING LOCAL CALIBRATION IN PRACTICE
We now focus on improving the local calibration of Neural Networks in practice. We first introduce Local Calibration Networks (LN), i.e., a two-component neural network architecture, designed to produce representations that exhibit improved local calibration.
Figure 1(a) shows the structure of LN. While standard post hoc calibrators consider the features representation as fixed, one component of LNs produces a new reduced-dimensionality feature representation $\phi^{\prime}(\mathbf{x}_{i})$ and the other component parametrizes new logits $\mathbf{l}^{\prime}$. From this perspective LNs can be viewed as a representation-level post-hoc calibration method. The two components are trained to minimize the misalignment between the probabilistic outputs and the local estimates of the class distribution computed from $\phi^{\prime}(\mathbf{x}_{i})$. To obtain these estimates, one can resort to kernel-based methods, e.g., Nadaraya-Watson estimators (Nadaraya, 1964; Watson, 1964), which require specifying a bandwidth hyper-parameter $\gamma$. More precisely, LNs minimize the following loss:
$\mathcal{L}_{\text{LN}}(\mathbf{x}_{i},\mathbf{y}_{i})=\frac{1}{n}\sum_{i=1}^{n}\overbrace{\text{d}_{\mathrm{JSD}}\left(\mathbf{\hat{p}}_{i},\hat{\theta}(\mathbf{y}_{i}\mid\phi^{\prime}(\mathbf{x}_{i})\right)}^{\text{Alignment term}}+$
$\lambda\cdot\underbrace{\mathcal{L}_{\text{ce}}\left(\mathbf{y}_{i},\hat{\theta}(\mathbf{y}_{i}\mid\phi^{\prime}(\mathbf{x}_{i})\right)}_{\text{Similarity term}},$ (9)
@@
On the one hand, the alignment term leverages the Jensen-Shannon distance between the model’s predicted probability vector and the kernel estimates for each instance in the training batch. We prove that, for a consistent estimator, the alignment term asymptotically converges to the divergence between the model prediction distribution and the true label distribution. This result relies on kernel consistency assumptions described in detail in Appendix A.
@@
Let $f$ be a probabilistic classifier and let $\hat{\theta}(\cdot\mid\mathbf{x})$ be a kernel estimator of $p(\cdot\mid\mathbf{x})$ that is consistent in mean squared error. Then, under standard regularity conditions:
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}\text{d}_{\mathrm{JSD}}\left(\mathbf{\hat{p}}_{i},\hat{\theta}(\mathbf{y}\mid\mathbf{x}_{i}\right)=$
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}\text{d}_{\mathrm{JSD}}\left(\mathbf{\hat{p}}_{i},\mathbf{p}_{i}\right).$ (10)
We provide proof in the Section B.5. Notably, the alignment term allows to learn probabilistic outputs that match the empirical neighbourhood frequencies.
On the other hand, the similarity term leverages the categorical cross entropy between the ground truth and the kernel estimates. This term serves a similar scope to the one of Lin et al. (2023). Intuitively, it encourages points with the same label to be attracted to nearby neighbors, but because the kernel is local, distant points of the same class exert little influence on each other. As a result, points that share fine-grained similarities are placed closer together, while more distinct variants remain further apart, but within the same class cluster. We provide a visual intuition of this behavior in Figure 1(b). However, the similarity term could still collapse class representations if not regularized or if the kernel bandwidth is too wide. Hence, we introduce another regularization hyper-parameter $\lambda$ (to be fine-tuned) to prevent this behaviour.
@@ 8 EXPERIMENTAL EVALUATION
  • Q1: Does our proposal outperform baselines on local calibration metrics?
  • Q2: Does our approach match the performance of baselines in global calibration metrics?
  • Q3: Does our method affect predictive performance?
The code [^2] Code developed by University of Trento authors. to reproduce our results can be found at https://github.com/Cesbar99/multiclass-local-calibration
8.1 Experimental Settings
@@
Method $ACC\,\uparrow$ $NLL\,\downarrow$
cifar10 cifar100 tissue cifar10 cifar100 tissue
LN $888\pm.002 .$$\underline{.688\pm.001}$ $630\pm.001 .$$347\pm.002$ $\mathbf{1.125\pm.002}$ $\mathbf{1.012\pm.003}$
DC $890\pm.002 .$$\mathbf{.690\pm.002}$ $617\pm.003 .$$\mathbf{.332\pm.007}$ $\underline{1.154\pm.009}$ $1.052\pm.009$
KO $\underline{.893\pm.001}$ $679\pm.003 .$$\underline{.632\pm.001}$ $340\pm.002$ $1.352\pm.009$ $1.027\pm.002$
KC $\mathbf{.893\pm.001}$ $679\pm.004 .$$\mathbf{.632\pm.001}$ $\underline{.340\pm.003}$ $1.351\pm.007$ $\underline{1.021\pm.002}$
PS $884\pm.001 .$$670\pm.002 .$$603\pm.008 .$$466\pm.004$ $618\pm.007 1.1$$180\pm.008$
IR $884\pm.001 .$$670\pm.002 .$$603\pm.008 .$$364\pm.008$ $1.437\pm.029$ $1.096\pm.016$
TS $884\pm.001 .$$670\pm.002 .$$603\pm.008 .$$362\pm.008$ $277\pm.009 1.1$$112\pm.023$
NC $884\pm.001 .$$670\pm.002 .$$603\pm.008 .$$494\pm.018$ $1.502\pm.036$ $2.100\pm.101$
Datasets. We evaluate our research questions over three multiclass datasets, i.e., cifar10, cifar100 (Krizhevsky et al., 2009) and tissuemnist from the MedMNIST collection (Yang et al., 2021).
Methods and Architectures. We evaluate our approach, LN, against different calibration baselines, including Temperature Scaling (TS(Guo et al., 2017), Isotonic Regression (IR(Zadrozny and Elkan, 2002), and Platt Scaling (PS(Platt and others, 1999). In addition, we compare with Dirichlet Calibration (DC(Kull et al., 2019), the current state-of-the-art for multiclass calibration. Moreover, we consider two K-Cal implementations (Lin et al., 2023): the first one considers the kernel classifier trained using the same loss as LN (KO); the second one does not use the extra Jensen-Shannon regularization as in the original version (KC). We also report the performance of the original non-calibrated classifier (NC).
All methods are based on a ResNet-50 backbone for CIFAR-10 and TissueMNIST, and a ResNet-152 for CIFAR-100. We detail hyperparameters in Appendix.
Metrics. For Q1, we compute two local calibration metrics, i.e., class-wise $LCE$ and $MLCE$. For Q2, we consider two global calibration metrics, i.e., $ECE$ and $ECCE$. For Q3, we report two performance measures, i.e., accuracy ($Acc$) and Negative Log-Likelihood ($NLL$). Implementation details are provided in Section D.3.
Experimental Setup. To evaluate calibration for all methods, we separate the data into three disjoint parts: $(i)$ a training set (further split into training and validation); $(ii)$ a calibration set (with an internal calibration/validation split); and $(iii)$ a held-out test set used exclusively for evaluation.
For CIFAR-10 and CIFAR-100, we use 45% of the original training data for training, 10% for validation, and 45% as test data, with the original test split serving as calibration data.
For TissueMNIST, we use the pre-computed splits, but since calibration metrics computation requires larger sample sizes to be meaningful, we split the training data in half, yielding balanced training and test splits ($\approx 82.5k$ each). The remaining pre-computed test set ($\approx 40k$ instances) is used for calibration.
For all the datasets, we split the calibration data into two sets, one to learn the calibration technique ($90\%$ of the calibration data) and one for validation (the remaining $10\%$ of the calibration data). Then, we $(i)$ train classifiers on the training set; $(ii)$ calibrate using the different methods on the calibration set; $(iii)$ compute the six metrics on the test set. We repeat this procedure using 5 different seeds and average results. Implementation details are provided in the Appendix.